A certificate of deposit mhas an annual rate of 6% compounded quarterly. Find the effective rate of interest. Show your work.
Answer by stanbon(57978)

A certificate of deposit has an annual rate of 6% compounded quarterly. Find the effective rate of interest.
---------------------------
A(t) = P(1+(r/n))^(nt)
--------------
Annual multiplier = (1+ (0.06)/4)^(4*1)
= (1+0.015)^4
= 1.06136
-------------
Annual interest rate = 0.06136 or 6.136%
==========================================
Cheers,
Stan H.